OpenStudy (anonymous):

solve this system in elimination 6x+5y=39 3x+5y=27

OpenStudy (displayerror):

Solve for either x or y with one equation, then substitute that into the other equation. So using the first equation, I'll solve for y:

OpenStudy (displayerror):

6x+5y = 39 5y = 39 - 6x y = (39 - 6x)/5 Now we know what y is equal to in terms of x. We can substitute that into the second equation, to get an equation that only has one variable, x. Now with that, you can solve for the numerical value of x, then plug that value into either equations 1 or 2 to solve for the other unknown, y.